OpenStudy (anonymous):

Determine whether the series is absolutely convergent, conditionally convergent or divergent. Give comprehensive reasons for your answer, using appropriate tests (see question for series)

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}(-1)^{n}\frac{ n }{5^{n}(n ^{2}+3)}\]

OpenStudy (anonymous):

i found it to be divergent using ratio test fo absolute divergence or convergence but i am doubting my answ

OpenStudy (anonymous):

\[\sum_{0}^{\infty} \left| (-1)^n \frac{n}{5^n (n^2+3)} \right| \le \sum_{0}^{\infty} \frac{1}{5^n} \le \sum_{0}^{\infty} \frac{1}{2^n} \le \infty\]Absolutely convergent by comparison test.

OpenStudy (anonymous):

damn!!! can u pliz take it step by step

OpenStudy (anonymous):

which func are you comparing the series to?

OpenStudy (anonymous):

I thought I did. If we take the absolute value, the (-1)^n term disappears. The fraction n/(n^2+3) is always less than one, so that means the sum of the absolute values have to be less than the sum of the 1/5^n terms. 1/5^n is always smaller than 1/2^n, which sums to 2, less than infinity implies absolutely convergent. How was that?

OpenStudy (anonymous):

and what is the sum?

OpenStudy (anonymous):

That wasn't part of the question.

OpenStudy (anonymous):

basically the function you are comparing it to is 1/5^n

OpenStudy (anonymous):

yeah, and 1/2^n

OpenStudy (anonymous):

and is it safe to say 1/5^n is a goemetric series with r=1/5 which is less than 1 hence convergent

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

but i have learnt that you consider the highest power in the numerator and denominator in finding your function to compare

OpenStudy (anonymous):

it doesn't seem right to spilt it like that

OpenStudy (anonymous):

The fraction n/(n^2+3) is always less than one, so we discarded it on the other side of the inequality sign.

OpenStudy (anonymous):

We have the sum of a bunch of terms. Each term has a few factors. When we take the absolute value, the (-1)^n factor is always one. The n/(n^2+3) factor is always less than one. Both of these are multiplied by the 1/5^n factor. That means the sum has to be less than the sum of the 1/5^n term alone.

OpenStudy (anonymous):

Thx i will go thru this stuff again till i get it

OpenStudy (anonymous):

Hang tough. Do math every day.

OpenStudy (anonymous):

i have had a look an instead of the comparison test i have rechecked my working and come up with convergence here it is below \[consider \lim_{n \rightarrow \infty}\left| (-1)^{n}\frac{ n }{5^{n}(n ^{2}+3} \right|\] =\[\frac{ 1 }{ 5 } \lim_{n \rightarrow \infty}\left| \frac{ n ^{3}+3n+n ^{2}+3 }{ n ^{3}+2n ^{2}+4n } \right|\] which becomes 1/5<1 hence convergent.L

OpenStudy (anonymous):

i skipped a few steps. does that look rite?

OpenStudy (anonymous):

No, recheck your work. The limit is zero.

OpenStudy (anonymous):

Here is my full working check the photo

OpenStudy (precal):

sorry, this is not my strong point

OpenStudy (anonymous):

thx,

hartnn (hartnn):

your work on limits seems correct to me, L=1/5.

OpenStudy (anonymous):

Thx @hartnn, this prob has given me a headache

hartnn (hartnn):

your welcome ^_^

OpenStudy (anonymous):

\[ \lim_{n \rightarrow \infty} \frac{n}{5^n(n^2+3)} \neq 1/5\] if you meant you used this term in the ratio test, where you divided the (n+1)th term by the nth term, then that would yield 1/5.

OpenStudy (anonymous):

You came in 3/4 way into a discussion follow the thread and you will notice that i was talking about using ratio test fo absolute divergence or convergence

OpenStudy (anonymous):

That's fine, but looking at your post you said \[ \lim_{n \rightarrow \infty} \left| (-1)^n \frac{n}{5^n(n^2+3)}\right| =\lim_{n \rightarrow \infty} (\text{blah blah blah}) = 1/5\] Which is not true, because the limit on the left is zero. Forgive my confusion, glad you got it worked out.